OpenStudy (anonymous):

(a^ 2n - a ^n - 6) (a^ n + 8) divide polynomials solve

OpenStudy (anonymous):

help me please

OpenStudy (anonymous):

Divide what by what and solve for what? I don't see a division symbol or an equals sign.

OpenStudy (anonymous):

(a^2n-a^n-6)/(a^n+8)

OpenStudy (anonymous):

So, find the quotient? How are you at long division?

OpenStudy (anonymous):



OpenStudy (anonymous):

I do not know but so is the question

OpenStudy (anonymous):

Ok, using the above picture and thinking in terms of long division, ask what do you have to multiply by a^n to get a^2n?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

do not know how you could help solve please

OpenStudy (anonymous):

Think about rules for exponents. When you multiply powers of the same base, you add the exponents; likewise, when you divide powers, subtract the exponents, so ... \[\large \frac{a^{2n}}{a^n} = a^{2n-n} = a^n.\] Does this make sense to you?

OpenStudy (anonymous):

a^n will be the first term in the quotient, so multiply that by the divisor and subtract to find the new dividend. What is a^n × (a^n +8)?

OpenStudy (anonymous):

okey thanks

OpenStudy (anonymous):

 This is your next step.